10. y – 2y + 2y = 2x, y(0) = 4, y0) = 8.​​​​​​

Find a solution
Answer #1

- equatich is, Here given given differential equation is yol-25+ 2y = 2x, y2O) = 4, so, Homogeneouas equation is, yl-2y+2y=0/ D doc =k do doc - for any constant k Yp(x) = AX+ 13 where A and 13 are constants. Then, yb (c) = dia K(1)-K and yocx) cu ((: ਹੈ Cuca).VCX)) uer) et oss + vek).cluor) de aer, d sinox) = cusci), alcosce sincx) and ar dx =CD) doc and dk=0 for any con